OpenStudy (anonymous):

Identify coordinates of the four vertices of (x+2)^2/9 + (y-3)^2/16 =1

OpenStudy (aum):

This is a vertical ellipse of the form: \( \Large \frac {(x-h)^2}{b^2} + \frac {(y-k)^2}{a^2} = 1 \) Compare the general equation above to the given equation and identify: a, b, h, k; where (h,k) is the center of the ellipse and 'a' is the semi-major axis and 'b' is the semi-minor axis. Knowing a, b, h, k, you can determine the coordinates of all four vertices.

OpenStudy (aum):



OpenStudy (anonymous):

So they would be 4,3,-2,3? @aum

OpenStudy (aum):

You have to say which is which. a = 4, b = 3, h = -2, k = 3 The center is (h,k) which is (-2,3). The top vertex will be at a height "a" ABOVE the center. That means, the y-coordinate will be 4 units above the center but the x-coordinate will be the same as the center. Top Vertex at (-2, 3+4) = (-2,7). Can you find the other three vertices?

OpenStudy (anonymous):

(-2,-1)?

OpenStudy (anonymous):

I'm not sure about the other 2

OpenStudy (aum):

Yes, the bottom vertex is (-2,-1). What is the left vertex? Notice how the left and right vertices are at the same height as the center which means their y-coordinates will be the same as that of the center. Only the x-coordinates will decrease by "b" units or "3 units" for the left vertex and increase by "3 units for the right vertex.

OpenStudy (anonymous):

oh, so they're (1,3) and (-5,3)?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

thank you :]

OpenStudy (aum):

You are welcome.